Assume that consumption is represented by C = 400 + 0.60Y. Also assume that investment (I) equals 200 and Xn equals 100. Calculate (a) the equilibrium level of income; (b) the multiplier; (c) consumption and saving at equilibrium level of income. If government spends additional 400, (d) what will be the new equilibrium income? e) What of consumption and saving? All figures are in billions of dollars. show all steps..
1
Expert's answer
2014-11-26T09:49:57-0500
C = 400 + 0.60Y, I = 200 and NX = 100.
(a) the equilibrium level of income can be found using the formula:
Y = C + I + G + NX
Y = 400 + 0.6Y + 200 + 100
0.4Y = 700
Y = $1750 billion
(b) the multiplier can be calculated using the formula:
m = 1/(1 - c) = 1/(1 - 0.6) = 2.5
(c) consumption and saving at equilibrium level of income will be:
C = 400 + 0.6*1750 = $1450 billion
S = s*Y = (1 - c)*Y = 0.4*1750 = $700 billion
If government spends additional 400, (d) the new equilibrium income will be 1750 + 400 = $2150 billion
e) Consumption will be: C = 400 + 0.6*2150 = $1690 billion
Saving will be: S = 0.4*2150 = $860 billion